OpenStudy (anonymous):

Whats the square root of 44c^3d^2

OpenStudy (anonymous):

\[\sqrt{44c ^{3}d ^{2}} = \sqrt{4*11*c*c*c*d*d} = 2cd \sqrt{11c}\]

OpenStudy (anonymous):

Square root of 45x^3y^3

OpenStudy (anonymous):

2cdSqrt11c